# LaCour!: Enabling Research on Argumentation in Hearings of the European Court of Human Rights

### v1

@@ 3.4 LaCour! statistics
On average, each hearing contains 14,038 word tokens and has an average sentence count of 557. With a standard deviation of 3,519 tokens, the majority of the transcripts are roughly of equal length. Exact statistics can also be found in Appendix A. Looking at the amount of tokens per hearing categorized by party outlined in Figure 4, one can see that there are several hearings with a large than average amount of tokens and very few hearing with a comparably smaller amount of tokens. If a third party is involved in the hearing, their part will usually be short. For applicant and government, the amount is also roughly equal. A closer look at the balance between applicant and government is also illustrated in Figure 4, which shows the absolute difference between the percentage of tokens spoken by the two opposing parties based on the length of the entire hearing. Apart from a few outliers, 85% of all hearings remain within a maximum of 20% difference between the two opposing parties. This means that generally, hearings have a balanced amount of time allocated for each party. The outliers can mostly be traced to the presence of more than one party on either side, which means that either there were multiple applicants or multiple respondents involved with fairly allocated speech time.
(a) Share of main languages based on time
(b) Duration of other languages in the share “Other”
@@
4 Preliminary experiments
In order to test and showcase our new corpus, we explore potential tasks which emerge with the new type of data for ECHR court proceedings. The LaCour! corpus partly bridges the gap between an application and a judgment, so we would like to investigate the impact of the hearing on the final outcome. While annotating the transcripts, we observed that some judges ask questions in almost every hearing, while others remain quiet. The questions often ask for some clarification on specific issues. The importance of questions in legal proceedings has also been studied by Johnson et al (2009) in the context of the U.S. Supreme Court. Their work, alongside our subjective observations during the annotation process, inspired us to take a closer look at the question-asking part of the hearings. Together with the questions asked during the hearings, the opinions give direct insight into an individual judge’s view on the case. We want to investigate whether there is a connection between a judge having a question during the hearing and a judge having a dissenting or concurring opinion in the subsequent judgment.
Judgments are formed by the Chambers and usually the majority of judges in the current Chamber have to agree on the verdict to form a majority opinion. There are multiple votes depending on the articles of the convention and aspects which are discussed. If an individual judge disagrees with the majority on a topic, they are asked to append their opinion in written form to the final judgment document. The most common types of judicial opinions are “dissenting”, which occur when a judge disagrees with the majority opinion and “concurring”, when a judge agrees with the outcome but for different reasons. In LaCour!, the linked judgment documents very frequently include these opinions. The most common type of opinion is “DISSENTING”, followed by “CONCURRING”. Some opinions are simply declared as “opinion” and some are partly dissenting and partly concurring. An opinion can be formulated by one or multiple judges. We extracted opinions of our linked documents with the help of regular expressions and further tagged them as “DISSENTING”, “CONCURRING”, “PARTLY” or simply “OPINION” if no further details are given. We excluded documents exclusively in French and opinions by judges who were not listed as present in the hearing and collected 758 opinions, 360 of which are unique. The distribution of the labels is shown in Table 1.
label unique opinions judge with opinion
DISSENTING 213 507
CONCURRING 94 182
OPINION 44 60
PARTLY 9 9
total 360 758
Table 1: Opinions with labels sourced from the judgment documents associated with the hearings; an opinion can be formulated by multiple judges
First, we conducted the exact Fisher test on all participating judges of our 154 hearings which are also listed in the judgment. We extracted matching questions from our transcripts by taking all content tagged with “Judge” [^18] This approach ignores potential questions asked by the Judge President, but it allowed us to extract questions without additional annotation or manual work and filtering questions with less than 25 tokens. This resulted in 651 questions (522 of them in English) with an average of 248 tokens. A cross tabulation of the extracted material, limited to only participants listed in both hearing and judgment, resulted in the numbers listed in Table 2.
has opinion
has question False True total
False 1304 530 1834
True 361 228 589
total 1665 758 2423
Table 2: Cross tabulation of participating judges in a hearing which do or do not have a questions in the hearing and an opinion in the associated judgment document
[…]
Based on these observations we inspected the connection between the content of question and opinion for the instances in which there was both a question in the hearing and an opinion in the judgment. We extracted this subset and limited it to only English questions, leaving us with 193 pairs. We then created sentence embeddings of each question and each opinion with a pre-trained Sentence-BERT (Reimers and Gurevych, 2019) model. [^19] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 The task of matching the questions to their opinions is formulated as a ranking task by finding the opinion embeddings with the highest cosine similarity for each question embedding. The pool of potential matches is limited to the amount of opinions we sourced (193). The precision@1 for this task is 0.16 and this result suggests that the content of question and opinion is not unrelated and should be investigated further.
As a more complicated task, we wanted to examine if it is possible to classify the type of the opinion based only on the question. A simple logistic regression or a multilayer perceptron (MLP) using the sentence embeddings as features was unable to learn anything from the set of questions labeled with the opinion tags shown in Table 1. In fact, we were unable to learn a classifier based on whether a question is followed-up with an opinion or not. However, since the opinion ranking task seems to suggest a connection on the basis of the sentence embedding, we further train a pre-trained Sentence-BERT model using contrastive loss and pairs of questions as an input. Although we were also unable to achieve results that outperform the majority baseline, we provide details for all our experimental setups in Appendix B.
Analysis and discussion
Our experimental results are preliminary to some extent. First, we used simple tools and models which can run with very low resource requirements. It is certainly possible that larger models are able to detect hints towards the ensuing opinion in the question. Second, our experiments do not consider the context of the questions, so future work might look into ingesting the remainder of the hearing or additional information. Overall, we show that our new LaCour! corpus allows for new kinds of interesting investigations that combine hearings with data from judgments.
@@ 5 Conclusion
We introduced our new corpus LaCour! which considers oral court proceedings in the ECHR for the first time. This corpus is available to the research community and enables the empirical analysis of ECHR court hearings. In addition to the transcribed and partly manually corrected text from the audio, we provided sentence-level timestamps and manually annotated role and language labels. Our collected data aims to reduce the gap that exists when considering ECHR proceedings as a whole, because it provides insights into the processes that takes place between the submission of an application and the final decision. The dataset characteristics can also enrich existing datasets or enable the formation of new datasets, which we showcased with our subset of questions and opinions. Although our preliminary experiments were limited to simple statistical tests and general sentence-embedding models, there are indications that there is possibly a connection between the questions asked in a hearing and opinions on a judgment. We believe that these experiments could be expanded on and our findings warrant further investigation. Apart from the use cases in legal NLP, we hope that law students or other interested parties will also use our data as a learning resource.
@@
There are some general limitations which originate from the material we worked with. One such limitation is that even though we try to illustrate most of the process of an application and case, we chose to use the original audio presented at the associated hearing. Although in our opinion, this is the best choice for understanding the speeches by applicants who did not speak English (or French), it should be noted that the judges and attendants of the hearings do not necessarily listen to the original audio. There is no way of knowing which interpretation a judge has heard, it could be either French or English. It is therefore not possible to draw conclusions from the choice of words used in the speeches and the judgments or opinions given by the judges of the Court. All names, institutions and parties were taken from the official documents provided in the HUDOC database. The only sections which are tagged with a name are those spoken by either “Judge” or “Judge President”. This enables linking questions to opinions in the judgment documents. However, these tags should not be used to automatically profile individual judges. We would also like to repeat that a large part of the transcript is the uncorrected output of a Whisper model. It is not unlikely that named entities and legal vocabulary are incorrectly transcribed. The quality of the transcriptions has not been evaluated for languages other than English and French. It should also be noted that under no circumstances do we wish to replicate or emulate a judge. We strongly believe that judgments concerning the lives of people should also be at the hand of an actual person.
@@ Acknowledgements
The independent research group TrustHLT is supported by the Hessian Ministry of Higher Education, Research, Science and the Arts. This work has been funded by the German Research Foundation as part of the ECALP project (HA 8018/2-1) and by the hessian.AI ConnectCom grant. We would also like to thank our research assistants who helped with the creation of the LaCour! corpus: Laura Boyette, Yassine Thlija, Marwa Naibkhil and Leonard Bongard.
    Appendix B Details on preliminary experiments

### v4

@@ 3.4 LaCour! statistics
On average, each hearing contains 14,038 word tokens and has an average sentence count of 557. With a standard deviation of 3,519 tokens, the majority of the transcripts are roughly of equal length. More descriptive statistics can also be found in Appendix A. Looking at the amount of tokens per hearing categorized by party outlined in Figure 4, one can see that there are several hearings with a larger than average amount of tokens and very few hearings with a comparably smaller amount of tokens. If a third party is involved in the hearing, their part will usually be short. For applicant and government, the amount is also roughly equal. A closer look at the balance between applicant and government is also illustrated in Figure 4, which shows the absolute difference between the percentage of tokens spoken by the two opposing parties based on the length of the entire hearing. Apart from a few outliers, 85% of all hearings remain within a maximum of 20% difference between the two opposing parties. This means that generally, hearings have a balanced amount of time allocated for each party. The outliers can mostly be traced to the presence of more than one party on either side, which means that either there were multiple applicants or multiple respondents involved with fairly allocated speech time.
@@
With the labels that we introduced to separate speakers, we can derive a hearing structure and separate parts of the hearing that were spoken before and after the break. The break is usually indicated by the announcer, although if this part is cut out of the video, it is still indicated by the judge president calling on the other judges for questions. With the labels we can also take a look at how the hearing structure has evolved over time. With the hearing structure, we can see that the number of questions asked by judges during hearings has increased over time. The trend can be observed in Figure 5. This is additionally supported by the percentage of tokens by judges in the question session of the hearing over time, shown in Figure 5, which is also increasing.
(a) Number of judges asking questions during the interactive part of the hearing over time
(b) Percentage of tokens spoken by judges in the interactive part of the hearing over time
Figure 5: Statistical overview of the hearing structure concerning the interaction and involvement of judge questions.
4 Experiments
LaCour! opens up many new possible points of investigation for ECHR proceedings. Our new corpus provides us with a window into the argumentation that was used before the judgment was formulated. Apart from the submissions by the parties, we also have insight into the interactions between judges and the parties. This motivates us to investigate the impact of a hearing on the final outcome.
4.1 Motivation and background
Besides the presentation of arguments by each party, which are prepared long before the hearing, there is a question section in the hearings. Judges may ask questions to the parties after the submissions and the parties prepare and present answers to these questions. The questions usually ask for some clarifications on specific issues. We have to assume that by the time the hearing takes place, the judges are familiar with the material and circumstances. The question session gives them the opportunity to ask for facts or reasoning by the parties and subliminally point out potential open issues or uncertainties. As indicated in the statistical analysis, the amount of questions has increased in the hearings over time, suggesting that it has become more important for the judges to interact with the parties.
Mathen et al (2024) thoroughly examined the importance and role of judicial opinions in the Canadian Supreme Court. They found that opinions play an important role and shape the value of a case outcome as a reference for future cases. Opinions are also used to mention and discuss key issues and points of ambiguity and disagreement on the interpretation of the law.
The importance of questions in legal proceedings has also been studied by Johnson et al (2009) in the context of the U.S. Supreme Court. This work also brings up the empirical question whether it is possible to detect a judge’s stance by their questions. The creation of linked corpora like Super-SCOTUS (Fang et al, 2023b) also enabled the empirical study of court hearings. Fang et al (2023a) observe that it is possible to detect linguistic clues to a judge’s stance in their questions during the oral arguments. They achieve an accuracy of 85% on the classification of partisanship by training a BERT (Devlin et al, 2018) model using only questions and text spoken by individual justices and the voting behavior.
This brings up the question we want to answer with our experiments. “Is a judge’s opinion on a case already made up by the time of the hearing and does the question give away hints to this opinion?” and “Is there a connection between a judge having a question during the hearings and a judge having a dissenting or concurring opinion in the subsequent judgment?” There is no partisanship in the ECHR, yet it is possible that we can detect clues that hint at disagreement with the majority.
4.2 Experimental setup
We approach the experiments in a similar way to (Fang et al, 2023a), by using only the questions by the judges as input and testing if the question alone gives away hints towards a judge’s position. For labels, we consider different configurations, the easiest being the type of opinion that is formulated by the same judge later. We extract all questions from the hearings, which we can easily select by considering all text tagged with the role label “Judge”. To connect the questions with the potential opinions attached to the judgment, we first link the hearing to the correct judgment file using the webcast-id and the type of the hearing (Chamber or Grand Chamber). From the connected judgment file, we extract all opinions attached after the actual judgment declaration. Using the opinion title, which contains the names of the judges who have written the opinion, we can extract the names as labels with regular expressions as well as the type of opinion (“dissenting”, “concurring”, “partly”, “opinion”). We limit the subset to only consider questions in English. After removing the name labels for anonymity, we end up with a sub dataset [^20] The dataset is available on https://huggingface.co/datasets/TrustHLT/QandO containing 475 individual questions linked to a matching opinion (or none).
We train various models from the BERT (Devlin et al, 2018) family to classify the questions by themselves with the matching label. We chose the basic BERT model to compare it to the results shown in  (Fang et al, 2023a), Legal-BERT (Chalkidis et al, 2020), assuming that it will have a decent performance due to the pretraining on texts from the legal domain, Legal-RoBERTa (Chalkidis et al, 2023) for the same reason and RoBERTa (Liu et al, 2019). For the RoBERTa models, we chose the large variant, to see if the model size affects the performance. Additionally, we train a classifier based on the state-of-the-art large language model Llama-3 8b (AI@Meta, 2024) using 4bit quantization with QLoRA (Dettmers et al, 2023). The number of instances with the label “partly” and “opinion” are severely underrepresented, so we remove them. The remaining labels are “dissenting”, “concurring” and “none” with 123, 50 and 282 instances respectively. For the final dataset with 455 instances, we use 70% for training, 15% for validation and reserve the remaining 15% for testing. Due to the imbalanced dataset, we preserve the distribution in the splits and use additional class weights in the training process to avoid overfitting. After the first round of experiments, we reduce the labels further to only classify in a binary manner if there is an opinion or not.
4.3 Results
The training details are shown in Appendix B and the results can be seen in Table 1. In the first setting, all models produce similar results, surpassing the random baseline, but not reaching a good stability as indicated by the high standard deviation. Especially the larger models struggle and none of the tested models produce decent results. Looking at the confusion matrix reveals that none of the instances from the class “concurring” were classified correctly. Surprisingly, Llama-3 8b does not manage to achieve scores that could be considered better than random guessing. The second setting appears to improve the performance of the models slightly. All models except Llama-3 8b surpass the random baseline. For the smaller models, the domain specific Legal-BERT has an advantage over BERT. But this difference in performance is not observable for the RoBERTa models. The best model in this setting is RoBERTa with a low standard deviation and 65% F1-score. We will provide an interpretation of these results later in section 4.5.
Setting Model F1-score ($\pm$ std) Accuracy ($\pm$ std)
dissenting, concurring, none Random $0.35\pm 0.03$ $0.25\pm 0.01$
BERT (base) $0.44\pm 0.04$ $0.35\pm 0.03$
Legal-BERT (base) $\textbf{0.45}\pm 0.05$ $\textbf{0.35}\pm 0.09$
RoBERTa (large) $0.41\pm 0.04$ $0.33\pm 0.07$
Legal-RoBERTa (large) $0.41\pm 0.05$ $0.35\pm 0.05$
Llama-3 8b $0.32\pm 0.10$ $0.24\pm 0.10$
dissenting, none Random $0.41\pm 0.02$ $0.36\pm 0.03$
BERT (base) $0.56\pm 0.05$ $0.48\pm 0.05$
Legal-BERT (base) $0.62\pm 0.08$ $0.55\pm 0.05$
RoBERTa (large) $\textbf{0.65}\pm 0.01$ $\textbf{0.58}\pm 0.01$
Legal-RoBERTa (large) $0.63\pm 0.06$ $0.55\pm 0.08$
Llama-3 8b $0.51\pm 0.07$ $0.52\pm 0.07$
Table 1: Results comparing all models using the labels “dissenting”, “concurring”, “none” in the first setting and “dissenting” and “none” in the second setting. Accuracy refers to the balanced accuracy. Both metrics are averaged over 3 runs.
4.4 Question embeddings and topics
Complementary to the classification experiments, we take a closer look if we can obtain more insights into the differences between questions with and without a corresponding judge’s opinion later on. To study this, we train a sentence transformer model (Reimers and Gurevych, 2019) based on Legal-BERT to embed the question. As training data, we use triplets consisting of a question, a positive example and a negative example. In our case, a positive example is a different question from the same class, a negative example is a different question from a different class. The classes are assigned by the presence and type of opinion of the judge in the same case in which the question was asked: “none”, “dissenting” and “concurring”. The underlying idea is that questions that appear in the same hearing and are followed up with the same type of opinion should be closer to each other, while questions that appear in the same hearing and resulted in different opinions should be further away from each other. During training, each question in a triplet is moved closer to its positive example and at the same time further away from its negative example. A sentence embedding reflects this by having a higher cosine similarity with similar instances. We avoid duplicates by removing inverse pairs of questions and positive examples. This contrastive sampling generates a dataset of 792 triplets which we used to train a sentence embedding model. The samples have a distribution similar to the dataset used for classification with 442, 250 and 100 triplets for the labels “none”, “dissenting” and “concurring”. The resulting fine-tuned embedding model can classify triplets with a surprisingly high cosine accuracy of 97%.
We take a closer look by running a BERT-Topic (Grootendorst, 2022) model with these embeddings. To get some meaningful topics, we remove stops words and cluster the embeddings using UMAP (McInnes et al, 2018).
Figure 6: Clustered topics of all questions. Names are determined by the most frequent words in the cluster. The embeddings are represented on the 2 dimensions that they are reduced on for clustering.
The topics for all questions are displayed in Figure 6. When taking a look at the topics by class in Figure 7, we can observe that questions that have a dissenting opinion are more represented in some topics than they are for questions without an opinion. One such example would be the topic “children_life_child”, which appears almost as frequently in questions tagged with dissenting opinions as in questions tagged with no opinions, despite the higher amount of questions with no opinion. Although it would require an expert to carefully examine the content of the questions to confirm this observation, this could hint at potential topics which are more controversial than others. Instead of linguistic clues in the way a judge asks their question, the indicating factor could also be linked to specific topics. We believe that this area could be further investigated.
Figure 7: Selection of the 5 most frequent topics grouped by the opinion type
4.5 Analysis
Despite high expectations because of the promising results on the similarly formulated task in (Fang et al, 2023a), we could not find such an overwhelming evidence that the questions asked by judges during the hearing can give us foreshadowing to an opinion or voting behavior. Some models managed to surpass the random baseline, yet the training is not stable for all models. Llama-3 8b was even unable to learn anything from the training data. If there is in fact a hint in those questions, it is likely too subtle to be picked up by state-of-the-art language models. Simply distinguishing between “dissenting” and “none” instead of the different types of opinions improves the performance of the models. Perhaps the difference between a concurring and a dissenting opinion is too difficult to detect. Alternatively it is possible that the number of examples for questions that resulted in a concurring opinion is too low. We observe that the relatively low performance actually sheds a positive light on the oral hearings in the ECHR, because it suggests that the language used is not as biased as it is in the U.S. Supreme Court. Similar to what Fang et al (2023a) find, we can also see that a legal domain-specific model does not necessarily work better than a general model. Our best performing model RoBERTa managed to achieve 65% F1-score on the task of classifying a question that is followed up by a dissenting opinion. These results show that the questions might not be completely free of hints and foreshadowing after all.
Through our experiments using sentence embeddings to represent the question topics, we find that dissenting opinions have certain topics in which they are more represented. This suggests that there are topics which are more controversial in the ECHR and on which the judges disagree more often. Our experiments demonstrated that material which we can extract from the hearings can bear relevancy to the final outcome and we are able to draw a connection between them.
An idea to further investigate this could be the inclusion of audio or video material to detect emotionality. Another possibility would be the inclusion of more transcript material to put the question into context. Yet another approach, which would likely require additional labeled data, could isolate the key issues that the judge’s questions address and compare them to the arguments brought forth by both parties on this particular issue. It could be interesting to see which issues are highlighted by judges, how they are addressed by the parties and what the final view of the court is on the issue. A more fine-grained approach like this would dive deeper into the exact argumentation used in hearings. However, this goes beyond the questions that we attempted to answer with our experimental setup, so we leave this for future work.
@@ 5 Conclusion
We introduced our new corpus LaCour! which considers oral court proceedings in the ECHR for the first time. This corpus is available to the research community and enables the empirical analysis of ECHR court hearings. In addition to the transcribed and partly manually corrected text from the audio, we provided sentence-level timestamps and manually annotated role and language labels. Our collected data aims to partly close the gap that exists when considering ECHR proceedings as a whole, because it provides insights into the processes that takes place between the submission of an application and the final decision. The dataset characteristics can also enrich existing datasets or enable the formation of new datasets, which we showcased with our subset of questions and opinions. LaCour! will also enable the study of argumentation in hearings, because it contains both the parties’ submissions and replies, as well as the questions posed by judges. This difference in modality and style of legal argumentation will also open up the possibility to study interesting novel questions, such as which arguments or views end up in the final judgment.
Although our experiments were limited to using only the questions from the interactive part of the hearing, there are indications that there is possibly a connection between the questions asked in a hearing and opinions on a judgment. We believe that these experiments could be expanded on and our findings warrant further investigation. Apart from the use cases in legal NLP, we also hope that law students or other interested parties will use our data as an educational resource.
@@
There are some general limitations which originate from the material we worked with. One such limitation is that even though we try to illustrate most of the process of an application and case, we chose to use the original audio presented at the associated hearing. Although in our opinion, this is the best choice for understanding the speeches by applicants who did not speak English (or French), it should be noted that the judges and attendants of the hearings do not necessarily listen to the original audio. There is no way of knowing which interpretation a judge has heard, it could be either French or English. It is therefore not possible to draw conclusions from the choice of words used in the speeches and the judgments or opinions given by the judges of the Court. All names, institutions and parties were taken from the official documents provided in the HUDOC database. The only sections which are tagged with a name are those spoken by either “Judge” or “Judge President”. This enables linking questions to opinions in the judgment documents. However, these tags should not be used to automatically profile individual judges. We would also like to repeat that a large part of the transcript is the uncorrected output of a Whisper model. It is not unlikely that named entities and legal vocabulary are incorrectly transcribed. The quality of the transcriptions has not been evaluated for languages other than English and French.
Concerning the connection between a question and an opinion, there is a factor to be considered, which is the actual decision of violation or non-violation of an article. The composition of the vote is not provided by the court, so we can only assume that a dissenting opinion indicates a vote against the majority. While this could be used as a feature in our study, there is often more than a single vote and multiple articles are concerned in a case. Consequently, the opinion can also be very nuanced and multifaceted which requires a model that can take into account both the fine-grained decision and the content of the opinion. The topics presented in Figure 7 only represent the questions on a surface level and capture keywords used. Additionally, the exact keywords are not extracted in a fully deterministic manner, so variance is possible. It should also be noted that under no circumstances do we wish to replicate or emulate a judge. We strongly believe that judgments concerning the lives of people should also be at the hand of an actual person.
@@ Acknowledgements
This work has been supported by the German Research Foundation as part of the ECALP project (HA 8018/2-1), by the hessian.AI ConnectCom grant, and by the Research Center Trustworthy Data Science and Security (https://rc-trust.ai), one of the Research Alliance centers within the https://uaruhr.de. We would also like to thank our research assistants who helped with the creation of the LaCour! corpus: Laura Boyette, Yassine Thlija, Marwa Naibkhil, and Leonard Bongard.
Appendix A Hearing statistics
The hearings in Table 2 were eventually excluded from the corpus:
WebcastID Title Date Reason
233009_07112012 Sindicatul ‘Pastorul cel Bun’ v. Romania (no. 2330/09) 07.11.2012 contains incorrect voice overlay in audio
4221907_12112014 Gherghina v. Romania (no. 42219/07) 12.11.2014 declared inadmissible (non-exhaustion of domestic remedies)
7153714_11012017 Harkins v. the United Kingdom (no. 71537/14) 11.01.2017 Interim measure were lifted, application was deemed inadmissible
5842813_22112017 Berlusconi v. Italy (no. 58428/13) 22.11.2017 case withdrawn by applicant
6312915_13112018 Elçi v. Turkey (no. 63129/15), Ahmet Tunç and Others v. Turkey (no. 4133/16) and Tunç and Yerbasan v. Turkey (no. 31542/16) 13.11.2018 All applications were deemed inadmissible
359918_24042019 M.N. and Others v. Belgium (no. 3599/18) 24.04.2019 application deemed inadmissible
5415516_12062019 Slovenia v. Croatia (no. 54155/16) 12.06.2019 no jurisdiction to take cognisance of the application
2095814_11092019 Ukraine v. Russia (re Crimea) (no. 20958/14) 11.09.2019 decision partly admissible, but another hearing is planned for 2023
Table 2: Hearings which were excluded from the corpus and the reason for the exclusion
More detailed statistics on LaCour! can be seen in Table 3.
number of transcripts 154
date range 01.01.2012 - 31.12.2021
tokens (nltk) 2,161,833
range of tokens per transcript 6458 - 25742
average tokens per transcript (±$\pm\sigma$) 14,038 $\pm$ 3519
median tokens per transcript 15,163
quartile of tokens [25%, 50%, 75%] [12020, 13784, 15295]
vocabulary size 77,156
average sentences per transcript (±$\pm\sigma$) 557 $\pm$ 155.90
quartile of sentences [25%, 50%, 75%] [454, 550, 625]
average tokens per sentence (±$\pm\sigma$) 25.2 $\pm$ 19.50
average conversation turns per transcript (±$\pm\sigma$) 26 $\pm$ 7.68
quantile (0.85) token % Applicant > Government 17.98
quantile (0.85) token % Government < Applicant 21.35
Table 3: Additional descriptive statistics for the LaCour! corpus
Appendix B Experimental setup details